# Dynamic Strategy Adaptation in Multi-Agent Environments with Large Language Models

### v2

@@ 3 Simulation Results and Analysis
Figure 5: Qualitative comparison in Overcooked-AI. Left: PPO+LLM agents coordinate by specializing in distinct roles and maintaining clear handoffs. Right: PPO-only agents often block each other, duplicate roles, or stall at delivery. The frozen LLM encourages efficient collaboration by rewarding high-level intent and coordination.
3.1 Returns per Episode
@@
Figure 6: Burger Kitchen: Coordination Under Order Timing Noise. Left: PPO+LLM agents adapt to temporal perturbations and successfully coordinate task flow despite misaligned order signals. The tomato icon represents an off-timing ingredient request, which the symbolic feedback loop helps agents recover from. Right: PPO baseline agents fail to deliver the completed burger due to miscoordination, resulting in idle output and a missed delivery. This figure illustrates how symbolic feedback enables robust strategy adaptation under timing noise.
Figure 3 summarizes average returns per episode for PPO and PPO+LLM. In the Clean setting, where agents observe the environment without any disturbance, PPO fails to achieve any reward. This shows that the agent fails to develop cooperative strategies without step-by-step feedback guiding its actions during the episode. However, PPO+LLM consistently has a return of 9.8 with zero variance, demonstrating that binary LLM evaluations are sufficient to support reliable task completion.
@@
Under the Visibility Noise condition, where the agent’s input state is interrupted with a random amount of perceptual noise, PPO again collapses to a return of 0. This outcome shows PPO’s sensitivity to low-level visual disturbances. In contrast, PPO+LLM stays strong and achieves a perfect return of 9.8 again. Since it learns from prompt-driven evaluations that do not rely on precise raw data, PPO+LLM remains stable even when perception is inconsistent.
The shaping mechanism becomes even more impactful under task-level disruption. In the Order Timing Noise setting, where task arrivals are randomly delayed or reordered, PPO performs poorly and has an average return of 40.4-40.4- 40.4 (standard deviation 3.7). This exposes a lack of stable learning which leads to low-quality actions. PPO+LLM significantly mitigates this degradation, recovering over 10 points (mean: 30.3-30.3- 30.3, std: 4.4). By consistently evaluating overall task progression (e.g., timestep and pending orders), the LLM provides structure that mitigates the unstable time dynamics that PPO alone fails to handle.
This pattern stays the same in the Combined Noise condition, where both visibility and timing are altered. PPO deteriorates to a mean of 39.9-39.9- 39.9 (std: 3.6), while PPO+LLM recovers nearly 25% of that loss with a mean return of 29.6-29.6- 29.6 (std: 4.3). This shows how symbolic feedback helps learning when rewards are weak or inconsistent, by giving the agent a clearer understanding of progress.
@@
To illustrate how the system works in practice, we present a qualitative example from the Overcooked-AI environment (Figure 5). In this episode, one agent retrieves ingredients while the other prepares plates and manages deliveries. The frozen LLM rewards states where agents remain active, coordinate handoffs efficiently, and avoid interfering with one another. This leads to smooth, high-throughput cooperation. In contrast, the PPO-only baseline often results in role duplication, hesitation, or spatial conflicts between agents. This example shows that symbolic feedback improves not just numerical performance but also the structure and clarity of collaborative behavior.
Figure 8: Burger Kitchen: Coordination Under Combined Noise. Left: PPO + LLM agents handle both mis-timed ingredient requests (clock icon) and sensor corruption (pixel-noise tile), reassigning roles on the fly and completing delivery. Right: PPO baseline agents break the task pipeline; perception noise triggers hesitation, timing jitter causes the burger to arrive too late, and the meal sits idle ($\times$) at the window. The comparison shows how prompt-based symbolic feedback enables recovery from multiple simultaneous disturbances.
@@
3.2 Robustness to Noise
Figure 4 shows the average return per episode under task-level noise. We evaluate each method’s sensitivity to structured disturbances to determine its robustness. To simulate insufficient state visibility, visibility noise modifies the grid-level tensor prior to it reaching the policy network. PPO collapses under noisy conditions due to its dependence on unal tered input encodings for fine-grained value learning, resulting in no reward accumulation. In contrast, PPO+LLM maintains positive returns across both conditions. This demonstrates that symbolic evaluations from the LLM offer resilience by reinforcing cooperative intent, even when perception is corrupted or incomplete.
@@ 3.2 Robustness to Noise
Figure 9: Latency per decision step across environments. PPO+LLM adds modest computational cost but stays within real-time limits across all conditions.
3.3 Latency Tradeoff
@@
Figure 10: Task-Completion Rate (9.4$\geq$ 9.4≥ 9.4) across environments. PPO+LLM reliably completes full tasks in clean and perceptually noisy settings, while PPO consistently fails to reach the success threshold.
3.4 Task Completion Rate
To evaluate whether agents successfully completed the full collaborative task, we introduced a completion-rate metric that counts the percentage of episodes where the total return exceeded a threshold of 9.4. This cutoff was chosen based on empirical observations, where PPO+LLM agents consistently reached scores near 9.8 in successful episodes. Setting the threshold slightly below 9.8 allows for small variations while still representing full task completion.
@@ 3.4 Task Completion Rate
Figure 11: Convergence to Nash equilibrium in a 2-agent Burger Kitchen task. PPO-LLM agents transition from initial miscoordination (Step 1) to stable cooperative behavior (Step 4), resulting in a mutually optimal joint policy where neither agent benefits from deviating.
@@
4 Conclusion and Future Work
In this paper, we have proposed a framework that integrates frozen large language models with reinforcement learning agents to support real-time strategy shaping in cooperative multi-agent environments. We have introduced symbolic feedback in the form of prompt-based binary evaluations, which refined low-level decision-making without requiring LLM fine-tuning or online optimization. We have evaluated the system across four increasingly noisy Overcooked-AI environments and found that PPO+LLM agents consistently outperformed PPO baselines. In high-noise conditions, we have achieved up to a 26% improvement in average return, while maintaining inference latency below 1.05 milliseconds per decision step. While our framework supports real-time shaping across multiple decision steps, it does not incorporate dialogue-based interactions, such as conversational feedback from the LLM or language-mediated coordination between agents, which may offer richer strategic reasoning and alignment. These forms of interactive feedback and inter-agent communication represent promising directions for future extension, potentially enabling more adaptive and socially aware multi-agent behavior.
Beyond these extensions, several additional research directions remain open. Our current system treats the LLM as a one-shot evaluator. However, enabling iterative and bidirectional interaction between agents and LLMs could enhance strategic grounding and help disambiguate complex decisions. Future frameworks may also incorporate agent-to-agent communication via natural language, supporting joint planning, intent signaling, and belief sharing. While our agents receive symbolic shaping within an episode, they currently do not persist or generalize feedback across episodes. Introducing memory modules or belief tracking could facilitate long-term strategy formation. Finally, although Overcooked-AI serves as a structured benchmark, future research should investigate more complex domains such as competitive negotiation, open-ended roleplay, or real-time robotics. These settings require symbolic reasoning, adaptive coordination, and language-guided shaping, and will further advance the development of autonomous, socially aware, and interpretable multi-agent systems.

### v3

@@ III Simulation Results and Analysis
Figure 5: Qualitative comparison in Overcooked-AI. Left: PPO+LLM agents coordinate by specializing in distinct roles and maintaining clear handoffs. Right: PPO-only agents often block each other, duplicate roles, or stall at delivery. The frozen LLM encourages efficient collaboration by rewarding high-level intent and coordination. Results averaged over n=200 evaluation episodes.
III-A Returns per Episode
@@
Figure 6: Burger Kitchen: Coordination Under Order Timing Noise. Left: PPO+LLM agents adapt to temporal perturbations and successfully coordinate task flow despite misaligned order signals. The tomato icon represents an off-timing ingredient request, which the symbolic feedback loop helps agents recover from. Right: PPO baseline agents fail to deliver the completed burger due to miscoordination, resulting in idle output and a missed delivery. This figure illustrates how symbolic feedback enables robust strategy adaptation under timing noise. Results averaged over n=200 evaluation episodes.
@@
Figure 3 summarizes mean rewards per episode for PPO and PPO+LLM. In the Clean setting, where agents observe the environment without any disturbance, PPO fails to achieve any reward. This indicates that the agent is unable to develop cooperative strategies without step-by-step feedback guiding its actions during the episode. In contrast, PPO+LLM consistently achieves a reward of 9.8 with zero variance, demonstrating that binary LLM evaluations are sufficient to support reliable task completion. Under the Visibility Noise condition, where the agent’s input state is interrupted with a random amount of perceptual noise, PPO again collapses to a return of 0. This outcome shows PPO’s sensitivity to low-level visual disturbances. In contrast, PPO+LLM stays strong and achieves a perfect return of 9.8 again. Since it learns from prompt-driven evaluations that do not rely on precise raw data, PPO+LLM remains stable even when perception is inconsistent.
Figure 4 shows the average return per episode under structured disturbances, meaning controlled and repeatable disruptions to either perception (observation noise) or task flow (order timing noise). In the Order Timing Noise condition, where task arrivals are randomly delayed or reordered, PPO performs poorly and has an average return of 40.4-40.4 (standard deviation 3.7). Negative returns occur because agents accumulate time-step penalties and failed delivery costs without successfully completing enough tasks to balance these losses, meaning policies waste time or act inefficiently such that penalties outweigh rewards. PPO+LLM significantly mitigates this degradation, recovering over 10 points (mean: 30.3-30.3, std: 4.4) by consistently evaluating overall task progression (e.g., timestep and pending orders), which reduces instability caused by noisy timing dynamics. Under Visibility Noise, where the grid-level tensor is partially corrupted before reaching the policy network, PPO again collapses due to its reliance on unaltered input encodings for fine-grained value learning, resulting in no return accumulation. In contrast, PPO+LLM maintains positive returns by relying on symbolic evaluations that remain stable even when perception is degraded. Together, these results demonstrate that PPO+LLM is substantially more resilient than PPO under both timing and observation disturbances, highlighting the importance of symbolic feedback for robust coordination in noisy multi-agent environments.
This pattern stays the same in the Combined Noise condition, where both visibility and timing are altered. PPO deteriorates to a mean of 39.9-39.9 (std: 3.6), while PPO+LLM recovers nearly 25% of that loss with a mean return of 29.6-29.6 (std: 4.3). This shows how symbolic feedback helps learning when rewards are weak or inconsistent, by giving the agent a clearer understanding of progress. To illustrate how the system works in practice, we present a qualitative example from the Overcooked-AI environment (Figure 5). In this episode, one agent retrieves ingredients while the other prepares plates and manages deliveries. The frozen LLM rewards states where agents remain active, coordinate handoffs efficiently, and avoid interfering with one another. This leads to smooth, high-throughput cooperation. In contrast, the PPO-only baseline often results in role duplication, hesitation, or spatial conflicts between agents. This example shows that symbolic feedback improves not just numerical performance but also the structure and clarity of collaborative behavior.
@@
Figure 8: Burger Kitchen: Coordination Under Combined Noise. Left: PPO+LLM agents handle both mis-timed ingredient requests (clock icon) and sensor corruption (pixel-noise tile), reassigning roles on the fly and completing delivery. Right: PPO baseline agents break the task pipeline; perception noise triggers hesitation, timing jitter causes the burger to arrive too late, and the meal sits idle ($\times$) at the window. This comparison shows how prompt-based symbolic feedback enables recovery from multiple simultaneous disturbances.
@@
III-B Robustness to Noise
@@
Figure 9: Latency per decision step across environments. PPO+LLM adds modest computational cost but stays within real-time limits across all conditions. Results averaged over n=200 evaluation episodes.
III-C Latency Tradeoff
@@
Figure 10: Task-Completion Rate ($\geq 9.4$) across environments. PPO+LLM reliably completes full tasks in clean and perceptually noisy settings, while PPO consistently fails to reach the success threshold. Results averaged over n=200 evaluation episodes.
III-D Task Completion Rate
To evaluate whether agents successfully completed the full collaborative task, we introduced a completion-rate metric that counts the percentage of episodes where the total return exceeded a threshold of 9.4. As shown in Figure 10, this cutoff was chosen based on empirical observations, where PPO+LLM agents consistently reached scores near 9.8 in successful episodes. Setting the threshold slightly below 9.8 allows for small variations while still representing full task completion.
@@ III-D Task Completion Rate
Figure 11: Convergence to Nash-like equilibrium in a 2-agent Burger Kitchen task. PPO-LLM agents transition from initial miscoordination (Step 1) to stable cooperative behavior (Step 4), resulting in a joint policy resembling equilibrium stability, where neither agent benefits from deviating unilaterally.
@@
IV Conclusion and Future Work
In this paper, we have proposed a framework that integrates frozen large language models with reinforcement learning agents to support real-time strategy shaping in cooperative multi-agent environments. We introduced symbolic feedback in the form of prompt-based binary evaluations, which refined low-level decision-making without requiring LLM fine-tuning or online optimization. We evaluated the system across four increasingly noisy Overcooked-AI environments and found that PPO+LLM agents consistently outperformed PPO baselines. In high-noise conditions, we achieved up to a 26% improvement in average return while maintaining inference latency below 1.05 milliseconds per decision step. More broadly, this framework demonstrates the potential of dynamic strategy refinement for cooperative decision-making in real-time multi-agent systems, where stability and adaptability are both critical.
While our results highlight strong empirical performance, several limitations remain. We do not provide a formal mathematical proof of convergence to equilibrium, and our findings are based on observed empirical dynamics. The framework also treats the LLM as a one-shot evaluator, without modeling longer-term memory or belief persistence. Furthermore, dialogue-based interactions, such as conversational feedback from the LLM or language-mediated coordination between agents, were not included in this study, though they may enable richer strategic reasoning and alignment. Addressing these limitations by incorporating formal proofs, memory-augmented evaluation, and interactive language feedback will form the basis of future extensions to this work.
Beyond these extensions, several research directions remain open. Enabling iterative and bidirectional interaction between agents and LLMs could enhance strategic grounding and help disambiguate complex decisions. Future frameworks may also incorporate agent-to-agent communication via natural language, supporting joint planning, intent signaling, and belief sharing. While our agents receive symbolic shaping within an episode, they currently do not persist or generalize feedback across episodes. Introducing memory modules or belief tracking could facilitate long-term strategy formation. Finally, future research should explore more complex environments such as competitive negotiation, open-ended roleplay, or real-time robotics, and investigate theoretical convergence guarantees. These directions will further advance the development of autonomous, socially aware, and interpretable multi-agent systems.
Funding
This work received no specific grant from any funding agency in the public, commercial, or not-for-profit sectors.
Conflict of Interest
The authors declare no conflict of interest.
Data and Code Availability
The data and code supporting the findings of this study are available from the corresponding author upon reasonable request.